# TimeDiT: General-purpose Diffusion Transformers for Time Series Foundation Model

### v1

@@ Abstract
With recent advances in building foundation models for texts and video data, there is a surge of interest in foundation models for time series. A family of models have been developed, utilizing a temporal auto-regressive generative Transformer architecture, whose effectiveness has been proven in Large Language Models. While the empirical results are promising, almost all existing time series foundation models have only been tested on well-curated “benchmark” datasets very similar to texts. However, real-world time series exhibit unique challenges, such as variable channel sizes across domains, missing values, and varying signal sampling intervals due to the multi-resolution nature of real-world data. Additionally, the uni-directional nature of temporally auto-regressive decoding limits the incorporation of domain knowledge, such as physical laws expressed as partial differential equations (PDEs). To address these challenges, we introduce the Time Diffusion Transformer (TimeDiT), a general foundation model for time series that employs a denoising diffusion paradigm instead of temporal auto-regressive generation. TimeDiT leverages the Transformer architecture to capture temporal dependencies and employs diffusion processes to generate high-quality candidate samples without imposing stringent assumptions on the target distribution via novel masking schemes and a channel alignment strategy. Furthermore, we propose a finetuning-free model editing strategy that allows the seamless integration of external knowledge during the sampling process without updating any model parameters. Extensive experiments conducted on a varity of tasks such as forecasting, imputation, and anomaly detection, demonstrate the effectiveness of TimeDiT.
@@ 1 Introduction
Time series analysis is pivotal in a diverse set of AI applications, such as natural science (Cuomo et al., 2022; Wang et al., 2020), social science (Zhang et al., 2022; Sharma et al., 2021),sustainability (Krenn and Buffoni, 2023),health (Kaushik et al., 2020; Kamra et al., 2021), etc. These applications root in diverse domains  (Li et al., 2018; Bi et al., 2023; Cao et al., 2023c; Zhang et al., 2021; Ye and Gao, 2022), leading to time series with various distributions (Wang et al., 2023) and a divers set of analysis tasks, such as forecasting (Nie et al., 2022; Jia et al., 2024), imputation (Tashiro et al., 2021), anomaly detection (Tuli et al., 2022), etc. Even though considerable progress has been made in developing specialized models optimized for specific scenarios and individual tasks, an open question remains: Can a single time series foundation model excel across domains? Recent initiatives have explored the possibility of universal time series models on zero-shot setting (Ansari et al., 2024; Liu et al., 2024; Gruver et al., 2024; Cao et al., 2023b), drawing inspiration from large pre-trained language models in natural language processing(NLP) and computer vision(CV), such as GPT(Radford et al., 2018), CLIP (Radford et al., 2021), which are known for their robust transfer learning capabilities. However, due to the fundamentally different semantics between text/images and time series data, the unique challenges of achieving a truly flexible and general-purpose time series model remain an open problem.
Recently, the emergence of LLMs like GPT-4 (OpenAI, 2023) and LLaMA (Touvron et al., 2023) suggests the potential for building time series foundation models enabling a general solution to handle multiple time series distributions. Previous attempts typically build upon the transformer backbone, which has achieved state-of-the-art performance on various time series tasks, particularly in modeling long-term dependencies. However, the tokenization of time series data for transformers is especially sensitive to variations in data sources and sampling rates. Previous tokenization approaches with different schemes including token patching (Das et al., 2023; Woo et al., 2024a); discretization tokens (Talukder et al., 2024) and tokens based on time series features (Yue et al., 2022; Ansari et al., 2024; Rasul et al., 2023) have either fragmented the global information or have been constructed in a manner that inherently loses important information. Most, if not all, of them employ a channel independence strategy (Nie et al., 2023) or focus solely on univariate time series. Channel independence strategy, though beneficial in certain contexts, often overlooks the complex inter-temporal and cross-feature dependencies in practical applications and thus presents an opportunity for optimization (Zeng et al., 2023).
Moreover, compared with texts and images, time series exhibit unique characteristics such as missing values (Kollovieh et al., 2023), irregular sampling (Kidger et al., 2020; Cao et al., 2023a; Zhang et al., 2022), multi-resolution (Niu et al., 2023, 2022), etc. To address these challenges, a foundation model for time series must be capable of demonstrating flexibility across different scales to handle diverse inputs with varying distributions. However, these unique natures and challenges are not covered by the popular well-curated benchmark datasets(Li et al., 2018; Zhou et al., 2021; Alexandrov et al., 2020). As a result, most existing works, which are developed and evaluated primarily on these datasets, may not fully address the complexities encountered in real-world time series applications. In addition, time series processes are often governed by underlying physical principles (Li et al., 2021; Meng et al., 2022; Xiao et al., 2023) and can be guided by domain-specific textual information (Jin et al., 2023; Sun et al., 2023). However, integrating these diverse sources of information into a unified model poses further challenges, as the model must effectively leverage the relevant physics context while adapting to the unique characteristics and distributions of each domain. Addressing each of these issues requires innovative approaches in data preprocessing, model architecture, and training strategies to create models that can seamlessly handle the diverse and complex nature of time series data.
Current approaches to time series modeling lack a unified framework for handling the aforementioned diverse and imperfect data inputs, frequently prioritizing performance on well-curated benchmarks over addressing real-world challenges. Diffusion models, such as DDPM (Ho et al., 2020), offer a promising solution by framing data generation as a series of conditional transformations, effectively recasting density estimation as sequential reconstruction. Unlike autoregressive methods that generate future tokens sequentially, diffusion models can directly produce high-quality samples through a reverse denoising process. This process can be analogized to solving partial differential equations (PDEs), allowing for the natural incorporation of physics-based knowledge. This capability, combined with transformers’ strength in capturing temporal dependencies, presents an opportunity to develop a more versatile and robust time series foundation model. Such a hybrid approach could effectively address the complexities of real-world data while maintaining the flexibility to adapt to various forecasting tasks and data conditions.
In this work, we introduce TimeDiT—a diffusion transformer-based foundation model equipped with a standardized training pipeline for different shapes of input time series and tailored for diverse distributions and downstream tasks. TimeDiT leverages the Transformer architecture’s inherent ability to capture temporal dependencies through its attention mechanisms, while also benefiting from the scalability that allows for increased model capacity crucial for complex time series tasks. By adopting a diffusion model approach, TimeDiT treats time series holistically, avoiding the error accumulation issues common in autoregressive solutions. The model incorporates a novel comprehensive mask mechanism that enables a single, unified foundation model to handle multiple tasks without additional modules or parameters. This design naturally addresses real-world challenges such as multi-resolution data and missing values. During the sampling stage, TimeDiT introduces an innovative strategy to incorporate physics knowledge as an energy-based prior, supported by theoretical guarantees. This approach guides the reverse diffusion process using physics-based constraints, including partial differential equations, resulting in generated samples that adhere to known physical laws and domain-specific requirements, thereby enhancing sample quality and model applicability across various scientific and engineering contexts.
TimeDiT’s performance is rigorously evaluated through an extensive experimental setup encompassing over 20 diverse datasets from domains including traffic, weather, finance, etc. The model is benchmarked against more than 25 open-source baselines, ranging from linear-based models to diffusion-based models, transformer-based models, and other forecasting foundation models. These comprehensive experiments cover multiple challenging time series tasks, including in-domain and zero-shot probabilistic forecasting, imputation, anomaly detection, and synthetic data generation. TimeDiT demonstrated state-of-the-art or highly competitive results across these tasks, showcasing its effectiveness and efficiency as a foundation model for various time series applications. Notably, TimeDiT achieved new state-of-the-art CRPSsum scores on the Electricity and Traffic datasets for probabilistic forecasting. In addition, the results on zero-shot experiments show that our model can be used as a foundation model even without fine-tuning, although fine-tuning may be necessary in some cases. Furthermore, TimeDiT’s scalability and adaptability are evident in its ability to incorporate external knowledge, such as physical constraints, during the sampling stage. This feature allows for the generation of samples that better conform to known physical laws and domain-specific requirements. This combination of state-of-the-art performance, adaptability across diverse tasks, scalability, and the ability to incorporate domain-specific knowledge positions TimeDiT as a powerful and versatile foundation model, capable of addressing a wide spectrum of time series challenges and opening new avenues for advanced time series analysis across various fields.
In summary, our contributions can be summarized as three unfolds:
  • We introduce TimeDiT, a novel diffusion transformer-based foundation model for time series analysis. By combining the strengths of diffusion models and transformers, our approach offers a flexible architecture adaptable to various downstream tasks. The model incorporates a comprehensive mask mechanism for reconstruction pretraining and task-specific fine-tuning, ensuring a standardized training pipeline capable of handling diverse input shapes and distributions.
  • Unlike autoregressive approaches, TimeDiT addresses real-world challenges in time series data by directly processing multivariate inputs and employing a denoising process to generate cohesive target time series. This method effectively handles issues such as missing values and multi-resolution data. In addition, TimeDiT can generate time series that adhere to known physical laws and domain-specific requirements, enhancing its applicability in scientific and engineering contexts.
  • Evaluated on over multiple datasets across different domains and tasks, TimeDiT achieves state-of-the-art or competitive results. It excels in probabilistic forecasting, imputation, anomaly detection, and data generation, showcasing its versatility as a foundation model in both in-domain and zero-shot settings.
@@ 2 Related Work
General Purpose Time Series Model
In the past decades, researchers have excelled in designing sophisticated models for specific time series analysis tasks (Zhang et al., 2024b; Fan et al., 2024; Cao et al., 2020, 2022). However, in recent years, the emergence of large language models has inspired the development of general-purpose time series models (Zerveas et al., 2021; Zhang et al., 2024a; Garza and Mergenthaler-Canseco, 2023) and the field of time series has seen tremendous exploration efforts towards foundation models. (Gruver et al., 2024) simply encoded time series as strings while (Jin et al., 2023) converted time series into language representations by alignment. (Cao et al., 2023b) and (Pan et al., 2024) further incorporated decomposition technique and prompt design and generalizes to unseen data and multimodal scenarios. (Rasul et al., 2023) worked towards foundation model from a probabilistic perspective but only considered univariate time series only which rarely appears in real-life. Additionally, many studies started to follow a two-stage training paradigm of pretraining and finetuning (Chang et al., 2023; Dong et al., 2024; Nie et al., 2022). However, these works mainly focused on the forecasting task only (Woo et al., 2024a; Das et al., 2023). (Zhou et al., 2023) first adapted GPT2 as a general-purpose time series analysis model and extended it to various time series tasks. (Talukder et al., 2024) leveraged VQVAE as a tokenizer for transformer to handle time series tasks and (Ansari et al., 2024) employed a scaling and quantization technique to embed time series. For more detailed literatures of the general-purpose time series model, please refer to recent surveys and position paper (Liang et al., 2024; Jin et al., 2024; Jiang et al., 2024)
@@
Despite the growing interest of diffusion models in various scenarios (Peebles and Xie, 2022; Li et al., 2022a; Lu et al., 2024; Sui et al., 2024a, b), the use of diffusions in time series analysis is less explored compared to pre-trained language models and transformers. Most existing studies also focused solely on forecasting and the choice of backbone model also varies among VAE(Li et al., 2022b), RNN(Rasul et al., 2021), and transformer. CSDI (Tashiro et al., 2021) utilized diffusion model for time series imputation. (Yuan and Qiao, 2024) incorporated decomposition into diffusion model to improve interoperability. Although  (Kollovieh et al., 2023) build a diffusion pipeline for multiple tasks with refinement, they still train different models for each task. To the best of our knowledge, there has been no exploration of leveraging unified diffusion models for a comprehensive set of time series tasks yet. Please refer to  (Yang et al., 2024) for a comprehensive literature review on diffusion models for time series analysis.
3 Preliminaries
3.1 Diffusion Model
In recent years, diffusion models have emerged as a promising approach in generative modeling. A diffusion process is a Markov chain that incrementally adds Gaussian noise to data over a sequence of steps, effectively destroying the data structure in forward process and destroying the data structure in backward structure.
By iterating this reverse process from $t=T$ down to $t=0$, the model gradually reconstructs the original data from noise. The reverse process learns to predict the mean and covariance of each intermediate distribution, effectively approximating the original data distribution.
@@ 4 Methodology
Figure 1: TimeDiT Architecture. Left: TimeDiT framework with diverse multivariate time series from different domains with multi-resolution, missing values; Middle: Structure of TimeDiT block; Right top: Illustration of masks generated by Time Series Mask Unit, reconstruction mask is neglected as it’s an all-zero mask; Right bottom: Masks for downstream tasks that TimeDiT handles during inference.
In this section, we present our main contributions: the proposed foundation model, TimeDiT , a diffusion model with transformer backbone designed for multiple time series tasks, along with uniform masking strategies and incorporation of physics knowledge and textual information as an extension. We first outline the uniformed problem setting for multiple down-stream tasks and offer an in-depth examination of the model architecture. Subsequently, we delve into the training pipeline with mask strategies, which help to build the training scheme in self-supervised learning for time series. Next, we present how to incorporate external information to improve the model’s performance during both the training and inference stages. By doing so, TimeDiT can generate samples that better conform to real-world requirements and enhance its performance on various downstream tasks. These extensions showcase the flexibility and adaptability of our proposed model, making it a powerful tool for a wide range of time series applications.
@@ 4.1 Problem Definition
We denote a multivariate time series as $\mathbf{X}=\{x_{i,j}\}\in\mathbb{R}^{K\times L}$, where $K$ is the number of features and $L$ is the length of the time series. Each individual entry $x_{i,j}$ represents the $j$-th feature at time step $l$, for $i\in\{1,\dots,K\}$ and $j\in\{1,\dots,L\}$. We define an observation mask $\mathbf{M_{obs}}=\{m_{i,j}\}\in\{0,1\}^{K\times L}$, where $m_{i,j}=0$ if $x_{i,j}$ is missing, and $m_{i,j}=1$ if $x_{i,j}$ is observed. Let $\mathbf{x}_{0}^{\text{obs}}\in X^{\text{obs}}$ denote the observed subsequence; $\mathbf{x}_{0}^{\text{tar}}$ denote the target subsequence of $\mathbf{x}_{0}^{obs}$ which could be forecast target or imputation target or the whole sequence depending on the task. Let $\mathbf{x}_{0}^{\text{con}}$ denote the unmasked partial observations in $\mathbf{x}_{0}^{obs}$ which acts like conditions for the masked area $\mathbf{x}_{0}^{tar}$. Let us use all subscripts of $x$ to denote diffusion timestamp, and a subscript of 0 means no noise has been applied to the original data. Formally, the goal of our task is to approximate the true conditional data distribution given the conditional information $q_{\mathbf{X}}\left(\mathbf{x}_{0}^{\mathrm{ta}}\mid\mathbf{x}_{0}^{\mathrm{con}}\right)$ with a model distribution $p_{\theta}(\mathbf{x}^{\text{tar}}_{0}\mid\mathbf{x}^{\text{con}}_{0})$, which can be calculated by a diffusion model with conditional information:
$p_{\theta}\left(\mathbf{x}_{0:T}^{\mathrm{tar}}\mid\mathbf{x}_{0}^{\mathrm{con}}\right):=p\left(\mathbf{x}_{T}^{\mathrm{tar}}\right)\prod_{t=1}^{T}p_{\theta}\left(\mathbf{x}_{t-1}^{\mathrm{tar}}\mid\mathbf{x}_{t}^{\mathrm{tar}},\mathbf{x}_{0}^{\mathrm{con}}\right),\mathbf{x}_{T}^{\mathrm{tar}}\sim\mathcal{N}(\mathbf{0},\mathbf{I}),\text{where}$
@@
The mask mechanism $\mathbf{M}$ plays a critical role in identifying the positions of $\mathbf{x}_{0}^{\mathrm{con}}$ and $\mathbf{x}_{0}^{\mathrm{tar}}$. By leveraging these positional differences, our model adeptly adapts to various downstream tasks, including forecasting, imputation, anomaly detection, etc, within a unified framework.

### v2

@@ Abstract
Foundation models, particularly Large Language Models (LLMs), have revolutionized text and video processing, yet time series data presents distinct challenges for such approaches due to domain-specific features such as missing values, multi-resolution characteristics, etc. Furthermore, the de-facto autoregressive transformers tend to learn deterministic temporal dependencies within pre-trained data while overlooking inherent uncertainties and lacking integration of physical constraints. In this paper, we introduce TimeDiT, a diffusion transformer model that synergistically combines transformer-based temporal dependency learning with diffusion-based probabilistic sampling. TimeDiT employs a unified masking mechanism to harmonize the training and inference process across diverse tasks while introducing a theoretically grounded, finetuning-free model editing strategy that enables flexible integration of external knowledge during sampling. Acknowledging the challenges of unifying multiple downstream tasks under a single model, our systematic evaluation demonstrates TimeDiT’s effectiveness both in fundamental tasks, i.e., forecasting and imputation, through zero-shot/fine-tuning; and in domain tasks, i.e., multi-resolution forecasting, anomaly detection, and data generation, establishing it as a proto-foundation model that bridges the gap between general-purpose and domain-specific models.
@@ 1 Introduction
Time series analysis is fundamental across natural science, sustainability, and healthcare (Ye et al., 2024; Kamra et al., 2021; Cuomo et al., 2022; Burger et al., 2024). While specialized models like TCNs (Franceschi et al., 2019), LSTMs (Siami-Namini et al., 2019), GNNs (Wu et al., 2020), and Transformers (Zhang and Yan, 2022) have advanced in time series tasks including forecasting, imputation, anomaly detection, and data generation, etc, their domain-specific design limits broader applicability. Inspired by the success of pre-trained models such as Deepseek Liu et al. (2024a), GPT-4 (OpenAI, 2023), LLaMA (Touvron et al., 2023) and Vision Transformer (Dosovitskiy et al., 2021)’s success in addressing foundation problems in language and vision, researchers have begun exploring universal time series forecasting models (Ansari et al., 2024; Liu et al., 2024c; Gruver et al., 2024). While these initial efforts have shown promise in temporal dependency forecasting, developing a foundation model that adapts to diverse time series tasks without task-specific customization remains challenging.
Recent advances in time series forecasting foundation models have established promising directions, while also revealing opportunities for further enhancement. Current approaches employ diverse tokenization strategies—such as TimeMoE (Shi et al., 2024) and TimesFM Das et al. (2023)’s patching, and Moirai Woo et al. (2024a)’s multiple patching—which effectively manage varying sequence lengths based on fixed patch length, yet present opportunities for improved generalization across dynamically shifting data characteristics. The channel independence strategies (Nie et al., 2023) adopted by models like Timer (Liu et al., 2024c) and Chronos Ansari et al. (2024) enable efficient model scaling, while suggesting potential for enhanced modeling of the intricate interplay between temporal patterns and cross-channel dependencies in real-world time series data (TSD). Moreover, conventional auto-regressive based models typically learn a deterministic, unique mapping relationship from historical data, limiting their ability to capture the inherent uncertainties and stochastic nature of TSD.
Moreover, real-world time series data poses fundamental challenges that disrupt inherent pattern learning in transformer models. Missing values (Kollovieh et al., 2023) break temporal continuity, multi-resolution sampling (Niu et al., 2023) distorts cross-variable relationships, and irregular temporal intervals (Cao et al., 2023a) compromise temporal dependency learning. However, current benchmark datasets (Li et al., 2018; Zhou et al., 2021; Alexandrov et al., 2020) often fail to reflect such real-world TSD’s complexities, potentially leading to models that underperform in practical applications. In addition, time series processes are often governed by underlying physical principles (Meng et al., 2022). Incorporating physics knowledge can further enhance model performance and interpretability, especially in data-scarce domains. Addressing these challenges demands innovative solutions across data preprocessing, model architecture design, and training strategies to create robust models that can effectively handle diverse TSD with varying temporal and feature characteristics.
While a universal foundation model for all time series tasks remains an aspirational goal, we propose a pragmatic yet powerful solution that bridges foundation models with task-specific adaptability, addressing both immediate challenges and future scalability. The solution is designed to espouse these three foundational principles: First, comprehensive pre-training through extensive data exposure establishes a robust foundation for temporal pattern recognition. This foundational capability directly addresses fundamental tasks with uniform definitions and minimal constraints, enabling forecasting and imputation through zero-shot inference or fine-tuning. Second, for tasks requiring in-domain training, such as anomaly detection where patterns inherently vary across datasets, adaptable pipelines are needed that maintain consistency without model architecture updating. Third, the foundation model can incorporate physical constraints and domain knowledge, enhancing performance while ensuring solutions remain grounded in real-world applicability.
Following the aforementioned foundational principles, in this work, we introduce TimeDiT and position it as a proto-foundation model designed to process practical TSD across domains, frequencies, and sampling patterns. As a diffusion transformer-based (Peebles and Xie, 2022) approach, TimeDiT combines the transformer architecture’s generalizability and expertise in capturing temporal dependencies with diffusion models’ capacity to explore diverse solutions within a broad prior space, enabling the direct generation of high-quality samples. TimeDiT incorporates comprehensive time series mask units, including position mask, stride mask and block mask, for both task-agnostic pre-training and task-specific inference that offer flexibility in handling varying input shapes and enable self-supervised learning (SSL). Furthermore, during the sampling stage, TimeDiT can incorporate physics knowledge, such as pre-defined partial differential equations (PDEs), as a theoretically grounded energy-based prior, supported by theoretical guarantees. This approach guides the reverse diffusion process using physics-based constraints, resulting in generated samples that adhere to known physical laws and domain-specific requirements, thereby enhancing sample quality and model applicability across various scientific and engineering contexts. This principled approach naturally leads to our comprehensive three-part experimental evaluation: First, zero-shot and fine-tuning on canonical tasks including forecasting and imputation from electricity, traffic, climate, and finance domains. Second, domain-adaptive tasks require prior knowledge of data characteristics—specifically multi-resolution forecasting, anomaly detection, and targeted data generation—including climate, health care, finance, energy, hydraulics, and science domains. Third, we further examine the model’s capacity for knowledge integration through physics-informed constraints and multimodal fusion. Notably, the results of zero-shot experiments show that our model can be used as a foundation model even without fine-tuning, although fine-tuning may be necessary in some cases. In addition, TimeDiT achieved new state-of-the-art (SOTA) performance in uncertainty quantification (UQ) across real-world datasets for probabilistic forecasting with missing values or multi-resolution. The main contribution of our work can be summarized as the following threefold:
  • We introduce TimeDiT, a proto-foundation model that integrates transformer-based temporal modeling with diffusion sampling capabilities. Enhanced by unified masking mechanisms, TimeDiT transcends conventional foundation models’ focus on forecasting, enabling comprehensive SSL for diverse time series tasks.
  • We develop a model-editing-free physics knowledge injection framework that leverages physical equations (such as PDEs) as energy-based priors. This approach ensures that generated samples adhere to known physical laws while maintaining model integrity.
  • We validate TimeDiT through comprehensive experiments aligned with our foundational principles, evaluating its performance across zero-shot, fine-tuning, and in-domain settings. The results demonstrate state-of-the-art performance, cross-task adaptability, and seamless integration of domain knowledge.
@@ 2 Related Work
Large Pre-trained Time Series Model
In the past decades, researchers have excelled in designing sophisticated models for specific time series analysis tasks (Zhang et al., 2024b; Fan et al., 2024; Cao et al., 2020; Bi et al., 2023; Zhang et al., 2021; Ye and Gao, 2022; Jia et al., 2024). However, the recent emergence of LLMs has inspired the development of general-purpose time series models and the field of time series has seen tremendous exploration efforts towards foundation models (Zerveas et al., 2021; Zhang et al., 2024a) . Specifically, (Gruver et al., 2024) simply encodes time series as strings while TimeLLM (Jin et al., 2023) convertes time series into language representations by alignment. TEMPO (Cao et al., 2023b) and S2IP-LLM (Pan et al., 2024) further incorporate decomposition technique and prompt design and generalize to unseen data and multimodal scenarios. Additionally, many studies start to follow a two-stage training paradigm of pretraining and finetuning (Chang et al., 2023; Dong et al., 2024). However, previous works including Timer (Liu et al., 2024c), TimeMoE Shi et al. (2024), Chronos (Ansari et al., 2024), TimeGPT (Garza and Mergenthaler-Canseco, 2023), UniTime (Liu et al., 2024b), TTM (Ekambaram et al., 2024) and Moirai (Woo et al., 2024b) mainly focus on the forecasting task only. (Zhou et al., 2023a) first adapted GPT2 as a general-purpose time series analysis model and extended it to various time series tasks. (Talukder et al., 2024) leveraged VQVAE as a tokenizer for transformer to handle time series tasks and (Ansari et al., 2024) employed a scaling and quantization technique to embed time series. For more detailed literatures of the general-purpose and foundation time series model, please refer to recent surveys (Liang et al., 2024; Jin et al., 2024; Jiang et al., 2024).
@@
Despite growing interest in diffusion models across various scenarios (Li et al., 2022a; Lu et al., 2024; Sui et al., 2024a, b), their application in time series analysis remains less explored compared to pre-trained language models. Most existing studies also focus solely on forecasting and the choice of backbone model also varies among VAE(Li et al., 2022b), RNN(Rasul et al., 2021), and transformers. Recently, CSDI (Tashiro et al., 2021) first utilizes a diffusion model for time series imputation with a self-supervised approach. SSSD (Alcaraz and Strodthoff, 2023) combines the structured state space model with the diffusion model for imputation. ImDiffusion (Chen et al., 2023) leverages diffusion models as time series imputers to achieve accurate anomaly detection. D3VAE (Li et al., 2022b) proposes a generative time series forecasting method on top of VAE equipped with the diffusion model. Meanwhile, DiffusionTS (Yuan and Qiao, 2024) incorporates decomposition into the diffusion model to improve interoperability. Although TSDiff (Kollovieh et al., 2023) build a diffusion pipeline for multiple tasks with refinement, they still train different models for each task. Based on our knowledge, no unified diffusion transformer model has yet been explored for a comprehensive set of time series tasks. For a thorough literature review on diffusion models in time series analysis, please refer to  (Yang et al., 2024).
@@ 3 Methodology
Figure 1: TimeDiT Architecture. Left: TimeDiT framework with diverse multivariate time series from different domains with multi-resolution or missing values; Middle: Structure of TimeDiT block; Right top: Illustration of masks generated by Time Series Mask Unit; Right bottom: Masks for downstream tasks that TimeDiT handles during inference.
In this section, we present our main contributions: the proposed foundation model, TimeDiT , a diffusion model with the transformer backbone designed for multiple time series tasks. We first outline the uniform problem setting for multiple downstream tasks and offer an in-depth examination of the model architecture. Subsequently, we delve into the training pipeline with mask strategies, which help to build the training scheme in self-supervised learning for time series. Next, we present how to incorporate external information to improve the model’s performance during inference stages by generating samples that better conform to real-world requirements. These extensions showcase the flexibility and adaptability of our proposed model, making it a powerful foundation model for a wide range of time series applications.
@@ 3.1 Problem Definition
We denote a multivariate time series as $\mathbf{X}=\{x_{i,j}\}\in\mathbb{R}^{K\times L}$, where $K$ is the number of features and $L$ is the length of the time series. Each individual entry $x_{i,j}$ represents the $j$-th feature at time step $i$, for $i\in\{1,\dots,L\}$ and $j\in\{1,\dots,K\}$. We define an observation mask $\mathbf{M_{obs}}=\{m_{i,j}\}\in\{0,1\}^{K\times L}$, where $m_{i,j}=0$ if $x_{i,j}$ is missing, otherwise, $m_{i,j}=1$. Let $\mathbf{x}_{0}^{\text{obs}}\in X^{\text{obs}}$ denote the observed subsequence; $\mathbf{x}_{0}^{\text{tar}}$ denote the target subsequence of $\mathbf{x}_{0}^{\text{obs}}$ which could be forecast target or imputation target or the whole sequence depending on the task. Let $\mathbf{x}_{0}^{\text{con}}$ denote the unmasked partial observations in $\mathbf{x}_{0}^{\text{obs}}$ which acts like self-conditions for the masked area $\mathbf{x}_{0}^{\text{tar}}$. Let us use all subscripts of $\mathbf{x}$ to denote diffusion timestamp, and a subscript of 0 means no noise has been applied to the original data. Formally, the goal of our task is to approximate the true conditional time series distribution given the conditional information $q_{\mathbf{X}}\left(\mathbf{x}_{0}^{\mathrm{tar}}\mid\mathbf{x}_{0}^{\mathrm{con}}\right)$ with a model distribution $p_{\theta}(\mathbf{x}^{\text{tar}}_{0}\mid\mathbf{x}^{\text{con}}_{0})$, which can be calculated by a diffusion model with conditional information:
$\textstyle p_{\theta}\left(\mathbf{x}_{0:T}^{\mathrm{tar}}\mid\mathbf{x}_{0}^{\mathrm{con}}\right):=p\left(\mathbf{x}_{T}^{\mathrm{tar}}\right)\prod_{t=1}^{T}p_{\theta}\left(\mathbf{x}_{t-1}^{\mathrm{tar}}\mid\mathbf{x}_{t}^{\mathrm{tar}},\mathbf{x}_{0}^{\mathrm{con}}\right),\mathbf{x}_{T}^{\mathrm{tar}}\sim\mathcal{N}(\mathbf{0},\mathbf{I}),\text{where}$
@@
The mask mechanism $\mathbf{M}$ plays a critical role in identifying the positions of $\mathbf{x}_{0}^{\mathrm{con}}$ and $\mathbf{x}_{0}^{\mathrm{tar}}$. By leveraging these positional differences, our model can adeptly adapt to tasks like forecasting, imputation, and anomaly detection in a unified framework.